OpenStudy (anonymous):

If I want to make 135. kg of NO, and the reaction proceeds in 78.0% yield, how many kilograms of NH3 must I start with in the presence of excess oxygen?

OpenStudy (anonymous):

I know the answer is 98.1 b/c of the answer key but i dont understand how to get that. I keep getting very large numbers

OpenStudy (abb0t):

Well, you want to start with a balanced equation. Do you have that?

OpenStudy (anonymous):

Yeah I have 4NH3 + 5O2 = 4NO

OpenStudy (abb0t):

You have 10 oxygen on the left, and 4 on the right.

OpenStudy (anonymous):

oh. Is my equation incorrect though before i start balancing it out

OpenStudy (abb0t):

Yes, you need to balance everything.

OpenStudy (anonymous):

but theres an H in the lest from NH3 and no H on the right on my product side

OpenStudy (abb0t):

And you're missing hydrogen. Even though you may get hydrogen molecules released as gas, you need to account for those in your equation.

OpenStudy (anonymous):

so is the product NO + H

OpenStudy (abb0t):

Yes.

OpenStudy (anonymous):

2 NH3 + O2 = 2 NO + 6 H

OpenStudy (abb0t):

That looks about right! Did you try and calculate your answer now and still getting large numbrs? Usually, that's where people make the mistake which gives them odd number values.

OpenStudy (anonymous):

yes. Im doing (1/0.78) * 78 = 135g = 173 kg. Then 173 kg/ 0.03 g = 5766 moles of NO

OpenStudy (anonymous):

The 5766 sounds too large. I was divinding the amoount of NO to be made in a 78% yield and got 173 kb but ti find moles it has to be in grams. So its 170000 g / 30 g = 5666 moles

OpenStudy (abb0t):

\[4 NH_3 + 7 O_2 \rightarrow 4 NO_2 + 6 H_2O \] I believe this is combustion.

OpenStudy (anonymous):

I was going to then multiply the moles i got by the ratio of (2 moles NO/ 2 moles NH3) but its still a large number

OpenStudy (anonymous):

Ok so if its combustion how will it change the problem now?

OpenStudy (abb0t):

I just wrote it down balanced for you. Hold on, I am going to explain it to you right now. Sorry, I am a bit rusty in my gen chem but I think I got it lol

OpenStudy (anonymous):

no problem! Im jsut glad ur helping

OpenStudy (anonymous):

becuase no matter what i was trying it wasnt working

OpenStudy (anonymous):

do u mind leaving step by step solution to what i should do. I have to go outside but once i com back i can look at it?

OpenStudy (abb0t):

According to the balanced equation 4 moles of NH3 would produce 4 moles of NO2 if the yield was 100% (which would be the theoretical yield). The % yield is only 0.78% meaning only 0.78% of 4 moles of NO2 (0.78 x 4 = _____ moles) will be produced. Now, find the amount of moles produced by NH3 and divide by NO2 moles and use dimensional analysis to figure out kilograms. I hope this makes sense. Lol. It made sense as I did it in my head, but i cant find my calculator. Lol

OpenStudy (anonymous):

Ok thank you! I'll look at it after i get home! Gotta go out for a project.